OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 3x^2 – 24x + 24 = 0

OpenStudy (anonymous):

getting pretty routine right ?

OpenStudy (anonymous):

lol yeahish..

OpenStudy (anonymous):

\[x^2-8x+8=0\\ x^2-8x=-8\\ (x-4)^2=-8+4^2\\ ...\]

OpenStudy (anonymous):

this time you got the \(...\) instead of the \(111\) typo

OpenStudy (anonymous):

i guess I'm not understanding how u get the formula...

OpenStudy (anonymous):

lets go slow and do it step by step

OpenStudy (anonymous):

lol alright...

OpenStudy (anonymous):

\[ 3x^2 – 24x + 24 = 0\] we need a leading coefficent of one, so divide all by \(3 \)

OpenStudy (anonymous):

ok... here hold on one sec.

OpenStudy (anonymous):

i am holding whilst you divide